# Inference-friendly Graph Compression for Graph Neural Networks

### v1

@@ 1. Introduction
(2) Real-time Inference. $\mathsf{GNN}$ models have been developed for e.g., traffic analysis (Peng et al., 2020), social recommendation (Fan et al., 2019), energy forecasting (Karimi et al., 2021; Fan et al., 2023; Ahmed et al., 2022), cybersecurity (Zhou et al., 2019), computer vision (Shi and Rajkumar, 2020), and edge devices (Zhou et al., 2021). Such scenarios often require real-time response at e.g., milliseconds  (Zhou et al., 2021). In such cases, even a linear time inference of “small” $\mathsf{GNNs}$ (when $F$ and $L$ are small constants) may still not be feasible for large graphs (when $|V|$ and $|E|$ are large).
(3) Fine-tuning & Benchmarking. Fine-tuning and testing pre-trained $\mathsf{GNNs}$ to adapt them for various domain-specific tasks is a routine process in $\mathsf{GNN}$-based data analysis for e.g., materials sciences, biomedicine, social science, and geosciences (Han et al., 2021; Zhili et al., 2024; Zhou et al., 2020; Reiser et al., 2022; Veličković, 2023). Inference tests of large pool of “candidate” $\mathsf{GNNs}$ over domain-specific graph data (such as knowledge graphs) is a cornerstone in such context. Fast $\mathsf{GNN}$ inference can accelerate large-scale domain-specific testing and benchmarking.
@@ 5.2. Anchored Graph Compression
6. Parallel Inference
$\mathsf{SPGC}$ and its variants go through the following modules and algorithms to parallelize inference, as illustrated in Fig. 7. (1) Coordinator. Given the configuration tuple $(M^{L},V_{T})$ provided by the user, coordinator conducts the following steps: (1) retrieves the corresponding $G$ and ${\mathcal{T}}$ from the compressor; (2) generates parallel joblets $<{\mathcal{T}}_{i},G_{c_{1}},M^{L}>$ and sends them to processor $P_{i}$ for all $v_{i}\in V_{T}$. (2) Compressor. Given a graph $G$ and a class of $\mathsf{GNNs}$ ${\mathcal{M}}^{L}$, the compressor produces the compressed graph $G_{c}$ and the memoization structure ${\mathcal{T}}$. It performs offline “once-for-all” graph compression by leveraging a built-in library of compression methods, including our structural-preserving compression $\mathsf{SPGC}$ as well as the state-of-the-art graph compression approaches (Liu et al., 2023; Kumar et al., 2023). (3) Parallel Inference. Given the joblet $<{\mathcal{T}}_{i},G_{c_{1}},M^{L}>$, the module computes the inference result $M^{L}(v_{i},G_{c_{i}})$ in parallel.
Coordinator. The coordinator optimizes query workload with three components. (1) Partitioner: Upon receiving an inference query $(M^{L},V_{T})$, it initializes a set of processors $\mathcal{P}$. At each $P_{i}\in\mathcal{P}$, it partitions $V_{T}$, the memoization table ${\mathcal{T}}$ to ${\mathcal{T}}_{i}$, and induces $G_{c_{i}}$ accordingly to $v_{i}\in V_{T}$. (2) Load Balancer: It assigns joblets to processors by distributing query workload evenly among the processors, preventing bottlenecks. This approach minimizes the total time cost of inference on the entire set of $V_{T}$. (3) Job Scheduler: It reschedules joblets $<\tau_{i},G_{c_{i}},M^{L}>$ based on estimated workload and current processor status, and allocate computational resources to dynamically rebalance the parallel inference computation.
Figure 7. Parallel Inference and Explainer with a running example: $M^{L}\in{\mathsf{GCN}}^{2}$; $V_{T}=\{a_{1},b_{1}\}$; ${\mathcal{T}}_{1}$, ${\mathcal{T}}_{2}$: partitioned subsets of ${\mathcal{T}}$ for $a_{1}$ and $b_{1}$; $G_{s_{1}}$, $G_{s_{2}}$: explanation graphs for $a_{1}$ and $b_{1}$.
Pipelining. At each processor $P_{i}$, given a joblet $<{\mathcal{T}}_{i},G_{c_{i}},M^{L}>$, it performs the inference for assigned $v_{i}\in V_{T}$ in parallel to generate output $M^{L}(v_{i},G_{c_{i}})$. The inference joblets follow a pipelined parallelism among all processes, to ensure that users receive inference result incrementally, rather than waiting or all inference queries to be evaluated.
Parallel Time Cost. The parallel cost per inference query is $O(\frac{L|G_{c}|F^{2}}{n})$ where $L$ is the number of layers and $F$ is number of features per node in $G_{c}$, and $n$ the number of processors.
@@ 7.1. Experimental Settings
Datasets. We employ four real-world graph benchmark datasets (summarized in Table. 5): (1) Cora (McCallum et al., 2000), a citation network where nodes represent documents, and edges are citations among the documents. Node features are described by a 0/1-valued word vector indicating the absence/presence of the corresponding word from the dictionary; (2) Arxiv (Hu et al., 2020), a citation network with nodes representing arXiv papers and edges denoting one paper cites another one. The features of each node includes a 128-dimensional feature vector obtained by averaging the embeddings of words in its title and abstract; (3) Yelp (Zeng et al., 2019) is prepared from the raw json data of businesses, comprising a dense network of user-business interactions. Nodes represent users and edges are generated if two users are friends. Node features are the word embeddings derived by reviews of products; and (4) Products (Hu et al., 2020), a product co-purchase network, where nodes represent products sold in Amazon and edges denote products purchased together. Node features are generated by extracting embeddings from the product descriptions followed by a Principal Component Analysis reducing the dimension to 100. The size of test nodes $|V_{T}|=0.05*|V|$ across all datasets. Note that $|V_{A}|=|V_{T}|$ since by default $V_{A}=V_{T}$ in $\mathsf{ASPGC}$.
Besides real-world benchmark datasets, we also generated a large synthetic dataset WS-MAG, by extending a core of real MAG240M network (Hu et al., 2021) (a citation network) with a small world generator  (Watts and Strogatz, 1998).
@@
Graph Neural Networks. We have pre-trained three classes of representative $\mathsf{GNNs}$: ${\mathsf{GCNs}}^{3}$ (Kipf and Welling, 2016), ${\mathsf{GATs}}^{3}$ (Velickovic et al., 2017), and ${\mathsf{GraphSAGE}}^{3}$ (Hamilton et al., 2017), for each dataset. For a fair comparison, (1) for all the datasets, we consider node classification, and (2) all compression methods are applied for the same set of $\mathsf{GNNs}$
@@ 7.2. ExperimentalResults
Varying Number of Layers. We varied the number of layers of $\mathsf{GNNs}$ from $2$ to $4$ over Arxiv and report its impact on inference speed-up (resp. accuracy) in Fig. 9(a) (resp. 9(b)). (1) $(0.5,1)$-$\mathsf{SPGC}$ consistently outperforms all the baselines in both inference speed-up and accuracy, due to that it preserves the inference results with small compressed graphs. (2) In general, the speed-up of inference achieved by $(0.5,1)$-$\mathsf{SPGC}$ is not sensitive to the number of layers. This verifies our theoretical analysis that it preserves inference results with unique smallest compressed graphs, which is independent of model complexity. (3) While the accuracy of $\mathsf{GNNs}$ drops as the number of layers become larger, in all cases, $(0.5,1)$-$\mathsf{SPGC}$ preserves the accuracy with smallest “gap” compared with other methods, for the same class of $\mathsf{GNNs}$.
@@
Varying $\alpha$. Fixing $r=1$, we varied $\alpha$ from $0.2$ to $1$, and report the results in Figs. 10(a) to  10(d). It tells us the followings.
@@
(3) As $\alpha$ increases, the F1-score (resp. accuracy) of the inference results over Yelp (resp. Products) remains insensitive, as shown in Figs.10(c) (resp. 10(d)). Our observation over Arxiv remains consistent, and we omitted it due to limited space. This indicates that $(\alpha,r)$-$\mathsf{SPGC}$ does not lose much on the quality of the inference while significantly improved inference efficiency.
@@
(c) $|G|$ v.s. Compression Cost
(d) Compression Cost
Figure 11. Compression: Scalability and Efficiency.
Impact of $|G|$. We increased the size of the simulated graph $|G|$ from $6M$ to $600M$, spanning two orders of magnitude. As shown in Fig. 11(c), (1) $\mathsf{SPGC}$, $\mathsf{ASPGC}$, and DPP all scale well with $|G|$, which is consistent with the compression cost presented in Table. 1. (2) $(\alpha,r)$-$\mathsf{SPGC}$ has relatively higher and more sensitive compression cost as $|G|$ increases, as $(\alpha,r)$-$\mathsf{SPGC}$ incurs more time to refine the $(\alpha,r)$-relation, yet achieves better compression ratio and in turn, more reduction of the inference cost on compressed graph $G_{c}$.
@@
Compression cost: real world datasets. Next, we compare the one-time compression cost induced by the $\mathsf{SPGC}$ or its variants with the cost of $\mathsf{DSpar}$ as shown in Fig. 11(d). Note that the compression cost of $\mathsf{SPGC}$ and its variants include the time cost of DPP. We have the following discoveries. (1) $\mathsf{SPGC}$, $(\alpha,r)$-$\mathsf{SPGC}$ and $\mathsf{ASPGC}$ outperforms $\mathsf{DSpar}$ on all the real-world datasets. In particular, $\mathsf{SPGC}$ are 95.52% and 58.45% times faster than $\mathsf{DSpar}$ over Arxiv and Products, respectively. For Yelp, $\mathsf{DSpar}$ does not run to completion after 1,000 seconds. (2) For $\mathsf{SPGC}$ and its variants, $\mathsf{ASPGC}$ outperforms $\mathsf{SPGC}$ and $(\alpha,r)$-$\mathsf{SPGC}$ by 12.71% and 68.34% on average. This is because $\mathsf{ASPGC}$ exploits data locality of $\mathsf{GNN}$ inference and perform compression up to certain hops of the anchored nodes
(a) Yelp
(b) Products
Figure 12. Total Cost: Varying Size of Queries
Total cost: Varying query workloads. We compare the total inference costs of $\mathsf{SPGC}$, $(\alpha,r)$-$\mathsf{SPGC}$, $\mathsf{ASPGC}$ to the total cost induced by the original $G$. For $\mathsf{SPGC}$ and its variants, the total cost is the sum of “one-time” compression cost and the inference cost for all inference queries; for “Original”, it refers to the inference time on the original graph. Varying the number of inference queries from $6$ to $18$ (each with fixed number of test node $|V_{T}|=0.05|V|$), Fig. 12 reports the total cost of a $3$-layers $\mathsf{GCN}$ on Yelp and Products.
[…]
@@
(2) Among $\mathsf{SPGC}$ variants, for queries with given anchored nodes, $\mathsf{ASPGC}$ performs best in improving the inference efficiency. We also observe that $(\alpha,r)$-$\mathsf{SPGC}$ is more sensitive to different datasets, compared with $\mathsf{SPGC}$ and $\mathsf{ASPGC}$. Indeed, $\mathsf{SPGC}$ enforces rigidly embedding equivalence, while $\mathsf{ASPGC}$ benefits most from data locality from anchored node set of fixed size, hence both are less sensitive to datasets. $(\alpha,r)$-$\mathsf{SPGC}$ on the other hand is most adaptive for tunable trade-off between inference speed up and compression ratio, depending on the heterogeneity of node embeddings.

### v2

@@ 1. Introduction
(2) Real-time Inference. $\mathsf{GNN}$ models have been developed for e.g., traffic analysis (Peng et al., 2020), social recommendation (Fan et al., 2019), forecasting (Karimi et al., 2021; Fan et al., 2023; Ahmed et al., 2022), computer vision (Shi and Rajkumar, 2020), and edge devices (Zhou et al., 2021). Such scenarios often require real-time response at e.g., milliseconds  (Zhou et al., 2021). In such cases, even a linear time inference of “small” $\mathsf{GNNs}$ (when $F$ and $L$ are small constants) may still not be feasible for large graphs (when $|V|$ and $|E|$ are large).
(3) Fine-tuning & Benchmarking. Fine-tuning and testing pre-trained $\mathsf{GNNs}$ to adapt them for various domain-specific tasks is a routine process in $\mathsf{GNN}$-based data analysis for e.g., materials sciences, biomedicine, social science, and geosciences (Han et al., 2021; Zhili et al., 2024; Reiser et al., 2022; Veličković, 2023). Inference tests of large pool of “candidate” $\mathsf{GNNs}$ over domain-specific graph data (such as knowledge graphs) is a cornerstone in such context. Fast $\mathsf{GNN}$ inference can accelerate large-scale domain-specific testing and benchmarking.
@@ 6.1. Experimental Settings
Datasets. We employ four real-world graph benchmark datasets (summarized in Table. 5): (1) Cora (McCallum et al., 2000), a citation network where nodes represent documents, and edges are citations among the documents. Node features are described by a 0/1-valued word vector indicating the absence/presence of the corresponding word from the dictionary; (2) Arxiv (Hu et al., 2020), a citation network with nodes representing arXiv papers and edges denoting one paper cites another one. The features of each node includes a 128-dimensional feature vector obtained by averaging the embeddings of words in its title and abstract; (3) Yelp (Zeng et al., 2019) is prepared from the raw json data of businesses, comprising a dense network of user-business interactions. Nodes represent users and edges are generated if two users are friends. Node features are the word embeddings derived by reviews of products; (4) Products (Hu et al., 2020), a product co-purchase network, where nodes represent products sold in Amazon and edges denote products purchased together; and (5) ogbn-papers100M, a billion-scale citation graph of 111 million papers and 1.6 billion edges. Its graph structure and node features are constructed in the same way as Arxiv. The size of test nodes $|V_{T}|=5\%*|V|$ across all datasets. For $\mathsf{ASPGC}$, $|V_{A}|=|V_{T}|$ as $V_{A}=V_{T}$ by default.
Besides real-world benchmark datasets, we also generated a large synthetic dataset WS-MAG, by extending a core of real MAG240M network (Hu et al., 2021) (a citation network) with a small world generator  (Watts and Strogatz, 1998). WS-MAG has a fixed number of nodes ($|V|$ = 2 million), while the number of edges $|E|$ increases linearly from 4 million to 1.192 billion.
@@
Graph Neural Networks. We have pre-trained three classes of representative $\mathsf{GNNs}$: ${\mathsf{GCNs}}^{3}$ (Kipf and Welling, 2016), ${\mathsf{GATs}}^{3}$ (Velickovic et al., 2017), and ${\mathsf{GraphSAGE}}^{3}$ (Hamilton et al., 2017), for each dataset. For a fair comparison, (1) for all the datasets, we consider node classification, and (2) all compression methods are applied for the same set of $\mathsf{GNNs}$.
@@ 6.2. ExperimentalResults
Varying Number of Layers. We varied the number of layers of $\mathsf{GNNs}$ from $2$ to $4$ over Arxiv and report its impact on inference speed-up (resp. accuracy) in Fig. 8(a) (resp. 8(b)). (1) $(0.5,1)$-$\mathsf{SPGC}$ consistently outperforms all the baselines in both inference speed-up and accuracy, due to that it preserves the inference results with small compressed graphs. (2) In general, the speed-up of inference achieved by $(0.5,1)$-$\mathsf{SPGC}$ is not sensitive to the number of layers. This verifies our theoretical analysis that it preserves inference results with unique smallest compressed graphs, which is independent of model complexity. (3) While the accuracy of $\mathsf{GNNs}$ drops as the number of layers become larger, in all cases, $(0.5,1)$-$\mathsf{SPGC}$ preserves the accuracy with smallest “gap” compared with other methods, for the same class of $\mathsf{GNNs}$.
@@
Varying $\alpha$. Fixing $r=1$, we varied $\alpha$ from $0.2$ to $1$, and report the results in Figs. 9(a) to  9(d). It tells us the followings.
@@
(3) As $\alpha$ increases, the F1-score (resp. accuracy) of the inference results over Yelp (resp. Products) remains insensitive, as shown in Figs.9(c) (resp. 9(d)). Our observation over Arxiv remains consistent, and we omitted it due to limited space. This indicates that $(\alpha,r)$-$\mathsf{SPGC}$ does not lose much on the quality of the inference while significantly improved inference efficiency.
@@
(c) Scalability: WS-MAG
(d) Scalability: ogbn-papers100M
Figure 10. Efficiency and Scalability at Billion scale.
Scalability Tests. We evaluated the scalability of $\mathsf{SPGC}$, $(\alpha,r)$-$\mathsf{SPGC}$ and $\mathsf{ASPGC}$ on large-scale graphs up to the billion-scale using WS-MAG (see Section 6.1) and ogbn-papers100M (Hu et al., 2020).
We varied the size of WS-MAG by enlarging its sizes with a generator that uses Watts-Strogatz algorithm (Watts and Strogatz, 1998). We varied the graph size $|G|$ from 6 million to 1.194 billion in a sequence. As shown in Fig. 10(c), We observe the following. (1) $\mathsf{SPGC}$, $\mathsf{ASPGC}$, and $(\alpha,r)$-$\mathsf{SPGC}$ scale well with $|G|$. This is consistent with our time cost analysis on the one-time compression in Table. 1. (2) $(\alpha,r)$-$\mathsf{SPGC}$ incurs a relatively higher and more sensitive compression cost due to the overhead of computing and refining the $(\alpha,r)$-relation. The overall cost remains to grow linearly with graph size, while its additional overhead yields improved compression ratios, leading to higher inference speed-ups on the compressed graphs $G_{c}$.
We also conducted a scalability test using a real-world billion-scale ogbn-papers100M (as illustrated in Fig. 10(d)). We generate subgraphs from 20% to 100% of the size of ogbn-papers100M using a subgraph sampling approach (Dupont et al., 2006) and evaluate the compression costs of $\mathsf{SPGC}$ and its variants. As the sampling rate increases, we observe that $\mathsf{SPGC}$, $\mathsf{ASPGC}$, and $(\alpha,r)$-$\mathsf{SPGC}$ all scale well. $(\alpha,r)$-$\mathsf{SPGC}$ incurs relatively higher and more sensitive compression costs, consistent with its performances on WS-MAG.
@@
Compression Cost Comparison. We compare the one-time compression cost induced by the $\mathsf{SPGC}$ or its variants with the cost of $\mathsf{DSpar}$ as shown in Fig. 11(a). Note that the compression cost of $\mathsf{SPGC}$ and its variants include the time cost of DPP. We have the following discoveries. (1) $\mathsf{SPGC}$, $(\alpha,r)$-$\mathsf{SPGC}$ and $\mathsf{ASPGC}$ outperforms $\mathsf{DSpar}$ on all the real-world datasets. In particular, $\mathsf{SPGC}$ are 95.52% and 58.45% times faster than $\mathsf{DSpar}$ over Arxiv and Products, respectively. For Yelp, $\mathsf{DSpar}$ does not run to completion after 1,000 seconds. (2) For $\mathsf{SPGC}$ and its variants, $\mathsf{ASPGC}$ outperforms $\mathsf{SPGC}$ and $(\alpha,r)$-$\mathsf{SPGC}$ by 12.71% and 68.34% on average. This is because $\mathsf{ASPGC}$ exploits data locality of $\mathsf{GNN}$ inference and perform compression up to certain hops of the anchored nodes.
(a) Compression Cost
(b) Total Cost
Figure 11. Time Cost Analysis (left: compression cost comparison; right: total cost for varying query loads).
@@
Total cost: Varying query workloads. We compare the total costs of $\mathsf{SPGC}$, $(\alpha,r)$-$\mathsf{SPGC}$, $\mathsf{ASPGC}$ to the total cost induced by the original $G$. For $\mathsf{SPGC}$ and its variants, the total cost is the sum of “one-time” compression cost and the inference cost for all inference queries; for “Original”, it refers to the inference time on the original graph. Varying the number of inference queries from $9$ to $18$ (each with fixed number of test node $|V_{T}|=0.05|V|$), Fig. 11(b) reports the total cost of a $3$-layers $\mathsf{GCN}$ on Products. (1) All $\mathsf{SPGC}$ methods are able to reduce the original inference cost, and all can improve the inference efficiency better with larger amount of queries. (2) Among $\mathsf{SPGC}$ variants, for queries with given anchored nodes, $\mathsf{ASPGC}$ performs best in improving the inference efficiency. We also observe that $(\alpha,r)$-$\mathsf{SPGC}$ is more sensitive to different datasets, compared with $\mathsf{SPGC}$ and $\mathsf{ASPGC}$. Indeed, $\mathsf{SPGC}$ enforces rigidly embedding equivalence, while $\mathsf{ASPGC}$ benefits most from data locality from anchored node set of fixed size, hence both are less sensitive to datasets. $(\alpha,r)$-$\mathsf{SPGC}$ on the other hand is most adaptive for tunable trade-off between inference speed up and compression ratio, depending on the heterogeneity of node embeddings.